### v1

# Thermal quasi-particle theory

Abstract
The widely used thermal Hartree–Fock (HF) theory is generalized to include the effect of electron correlation while maintaining its quasi-independent-particle framework. An electron-correlated internal energy (or grand potential) is defined by the second-order finite-temperature many-body perturbation theory (MBPT), which then dictates the corresponding thermal orbital (quasi-particle) energies in such a way that all thermodynamic relations are obeyed. The associated density matrix is of the one-electron type, whose diagonal elements take the form of the Fermi–Dirac distribution functions, when the grand potential is minimized. The formulas for the entropy and chemical potential are unchanged from those of Fermi–Dirac or thermal HF theory. The theory thus postulates a finite-temperature extension of the second-order Dyson self-energy of one-particle many-body Green’s function theory and can be viewed as a second-order, diagonal, frequency-independent, thermal inverse Dyson equation. At low temperature, the theory approaches finite-temperature MBPT of the same order, but it outperforms the latter at intermediate temperature by including additional electron-correlation effects through orbital energies. A physical meaning of these thermal orbital energies (including that of thermal HF orbital energies, which has been elusive) is proposed.
I Introduction
Thermal Hartree–Fock (HF) theoryHusimi (1940); Mermin (1963); Gu and Hirata (2024) is a curious ansatz. It uses an auxiliary one-electron Hamiltonian to define its one-electron density matrix, whereas a true density matrix for thermodynamics governed by an interacting Hamiltonian should be much higher ranked.Farid et al. (2000) Its diagonal elements, as variational parameters, ultimately become the Fermi–Dirac distribution functions of the non-interacting Fermi–Dirac theory, when the grand potential is minimized. The state energies defining this grand potential are, however, evaluated with the exact Hamiltonian containing two-electron interactions. Stipulated in this way, the theory does not seem to correspond to a single, consistent grand partition function.
Nevertheless, the thermodynamic functions of this theory are shownArgyres et al. (1974); Gu and Hirata (2024) to obey all thermodynamic relations. They also correctly reduceGu and Hirata (2024) to the zero-temperature HF theory.Szabo and Ostlund (1982); Shavitt and Bartlett (2009) They are widely used in applications to, e.g., energy bands in a solid where their temperature-dependent orbital energies are often invoked, somewhat unscrupulously, in explaining metal-insulator transitions,Hermes and Hirata (2015) etc. This is in spite of the facts that their physical meaning is still obscurePain (2011) and that state energies in quantum mechanics are supposed to be constant of temperature.
In this article, inspired by the immense success of thermal HF theory, we generalize this ansatz to include the effects of electron correlation, while keeping to its quasi-independent-particle framework and grand canonical ensemble. Our strategy of postulating what we call thermal quasi-particle theory is as follows:
We first assume that every state consists of non-interacting quasi-particles, each occupying a well-defined orbital having an electron-correlated orbital energy in a Slater determinant. The internal energy (or equivalently, the grand potential) is postulated by consulting with finite-temperature many-body perturbation theory (MBPT).Hirata and Jha (2019, 2020); Hirata (2021a) Its definition dictates the forms of the thermal quasi-particle energies (i.e., the electron-correlated thermal orbital energies) such that all thermodynamic relations are obeyed. The corresponding density matrix is then of the one-electron type, whose diagonal elements take the form of the Fermi–Dirac distribution functions, when the grand potential is minimized. The formulas for the entropy and average number of electrons (determining the chemical potential) are unchanged from the corresponding formulas of Fermi–Dirac or thermal HF theory. The thermal quasi-particle energies thus defined are interpretable as a finite-temperature generalizationMatsubara (1955); Luttinger and Ward (1960); March et al. (1967); Kadanoff and Baym (2018) of the Dyson self-energies of one-particle many-body Green’s function (MBGF) theory.Hirata et al. (2017)
Being based on perturbation theory, thermal quasi-particle theory forms a hierarchy of approximations with increasing accuracy and complexity. Fermi–Dirac and thermal HF theories can be viewed, respectively, as the zeroth- and first-order instances of this hierarchy. Here, we introduce a second-order thermal quasi-particle theory, which is the leading order in describing electron correlation, on the basis of the second-order grand potential or internal energy of finite-temperature MBPT.Hirata and Jha (2020); Hirata (2021a) We propose the expression for the corresponding second-order thermal self-energies, which obey all thermodynamic relations and ensure the variationality of the grand potential. The theory thus constitutes a second-order, diagonal, frequency-independent, thermal inverse Dyson equation.Hirata et al. (2017) It reduces to the second-order MBPT and MBGF (in the diagonal and frequency-independent approximation) at zero temperature.
A comparison of thermal quasi-particle theories with finite-temperature MBPT suggests that the former may include higher-order perturbation corrections through correlation-corrected thermal orbital energies and can thereby outperform the latter at intermediate temperatures. Furthermore, a comparison with thermal full-configuration-interaction (FCI) theoryKou and Hirata (2014) supports a proposed physical meaning of these thermal orbital energies, including that of thermal HF orbital energies, which has been elusive.Pain (2011); Hirata (2021a)
II Fermi–Dirac theory
Fermi–Dirac theory can be considered as the zeroth-order thermal quasi-particle theory [thermal QP(0) theory]. It is reviewed only briefly since it is discussed more fully in Ref. Gu and Hirata, 2024. Its internal energy $U^{(0)}$ and entropy $S^{(0)}$ at the inverse temperature $\beta=(k_{\text{B}}T)^{-1}$ are given by
$U^{(0)}$$\equiv$ $\langle E^{(0)}\rangle=\sum_{p}\epsilon_{p}^{(0)}f_{p}^{-},$ (1)
$S^{(0)}$$\equiv$ $-k_{\text{B}}\left(f_{p}^{-}\ln f_{p}^{-}+f_{p}^{+}\ln f_{p}^{+}\right)$ (2)
with the Fermi–Dirac distribution functions,
$f_{p}^{-}$ $=$ $\frac{1}{1+e^{\beta(\epsilon_{p}^{(0)}-\mu^{(0)})}},$ (3)
$f_{p}^{+}$ $=$ $1-f_{p}^{-}=\frac{e^{\beta(\epsilon_{p}^{(0)}-\mu^{(0)})}}{1+e^{\beta(\epsilon_{p}^{(0)}-\mu^{(0)})}}.$ (4)
Here, $\epsilon_{p}^{(0)}$ is the energy of the $p$th spinorbital of a reference wave function, and $\mu^{(0)}$ is the chemical potential determined by the electroneutrality condition,
$\bar{N}$ $=$ $\sum_{p}f_{p}^{-},$ (5)
where $\bar{N}$ is the average number of electrons, and the summations are always taken over all spinorbitals. The $\langle E^{(n)}\rangle$ in Eq. (1) is a zeroth-order thermal average of state energies, i.e.,
$\langle E^{(n)}\rangle$$\equiv$ $\frac{\sum_{I}E_{I}^{(n)}e^{-\beta(E_{I}^{(0)}-\mu^{(0)}N_{I})}}{\sum_{I}e^{-\beta(E_{I}^{(0)}-\mu^{(0)}N_{I})}},$ (6)
where $E_{I}^{(n)}$ is the $n$th-order perturbation correction to the energy of the $I$th Slater-determinant state according to Hirschfelder–Certain degenerate perturbation theoryHirschfelder and Certain (1974) with $E_{I}^{(0)}=\sum\epsilon_{i}^{(0)}$ ($i$ labels spinorbitals occupied in the $I$th state). Sum-over-states expressions of the thermal averages can be reduced to sum-over-orbital ones either by combinatorial logicHirata and Jha (2019, 2020) or by normal-ordered second quantization at finite temperature.Hirata (2021a)
The grand potential $\Omega^{(0)}$ is then given by
$\Omega^{(0)}$$\equiv$ $U^{(0)}-\mu^{(0)}\bar{N}-TS^{(0)}$ (8)
$=$ $\sum_{p}\left(\epsilon_{p}^{(0)}-\mu^{(0)}\right)f_{p}^{-}$
$+\frac{1}{\beta}\sum_{p}\left(f_{p}^{-}\ln f_{p}^{-}+f_{p}^{+}\ln f_{p}^{+}\right).$
As explicitly shown in Ref. Gu and Hirata, 2024, they satisfy the thermodynamic relations such as Eq. (8) and
$-\frac{\partial\Omega^{(0)}}{\partial\mu^{(0)}}$ $=$ $\bar{N},$ (9)
$-\frac{\partial\Omega^{(0)}}{\partial T}$ $=$ $S^{(0)}.$ (10)
The Fermi–Dirac distribution function [Eq. (3)] is a diagonal element of the one-electron density matrix that minimizes the grand potential. Therefore,
$\frac{\partial\Omega^{(0)}}{\partial f_{p}^{-}}$ $=$ $0,$ (11)
which is also easily verifiable.Gu and Hirata (2024)
Fermi–Dirac theory is an exact theory for thermodynamics of a system governed by an independent-particle Hamiltonian.
III Thermal Hartree–Fock theory
Thermal HF theoryHusimi (1940); Mermin (1963) is the first-order instance of thermal quasi-particle theory [thermal QP(1) theory], which is also fully discussed in Ref. Gu and Hirata, 2024. It accommodates the exact Hamiltonian with two-electron interactions within a quasi-independent-particle framework. Its internal energy $U^{\text{HF}}$ is postulated by an intuitively natural finite-temperature generalization of the zero-temperature HF energy,Hirata (2021a) i.e.,
$U^{\text{HF}}$$\equiv$ $\langle E^{(0)}\rangle+\langle E^{(1)}\rangle$ (12)
$=$ $\sum_{p}\epsilon_{p}^{\text{HF}}f_{p}^{-}-\frac{1}{2}\sum_{p,q}\langle pq||pq\rangle f_{p}^{-}f_{q}^{-}$ (13)
with
$\epsilon_{p}^{\text{HF}}$$\equiv$ $h_{pp}+\sum_{r}\langle pr||pr\rangle f_{r}^{-},$ (14)
where $h_{pq}$ is the one-electron (“core”) part of the Hamiltonian matrix element,Szabo and Ostlund (1982) $\langle pq||rs\rangle$ is an anti-symmetrized two-electron integral,Szabo and Ostlund (1982); Shavitt and Bartlett (2009) and $\langle E^{(1)}\rangle$ has been reduced in Eqs. (45) and (46) of Ref. Hirata and Jha, 2020. Spinorbitals labeled by $p$ and $q$ are the ones that bring the matrix of thermal HF orbital energies $\bm{\epsilon}^{\text{HF}}$ into a diagonal form, i.e.,
$\epsilon_{pq}^{\text{HF}}$ $=$ $h_{pq}+\sum_{r}\langle pr||qr\rangle f_{r}^{-}=\left\{\begin{array}[]{ll}\epsilon_{p}^{\text{HF}}&(p=q);\\ 0&(p\neq q),\end{array}\right.$ (17)
The Fermi–Dirac distribution functions,
$f_{p}^{-}$ $=$ $\frac{1}{1+e^{\beta(\epsilon_{p}^{\text{HF}}-\mu^{\text{HF}})}},$ (18)
$f_{p}^{+}$ $=$ $1-f_{p}^{-}=\frac{e^{\beta(\epsilon_{p}^{\text{HF}}-\mu^{\text{HF}})}}{1+e^{\beta(\epsilon_{p}^{\text{HF}}-\mu^{\text{HF}})}},$ (19)
are now defined with $\epsilon_{p}^{\text{HF}}$ and $\mu^{\text{HF}}$. The latter is determined by the same electroneutrality condition,
$\bar{N}$ $=$ $\sum_{p}f_{p}^{-}.$ (20)
The entropy formula is unchanged from that of Fermi–Dirac theory [Eq. (2)],
$S^{\text{HF}}$$\equiv$ $-k_{\text{B}}\sum_{p}\left(f_{p}^{-}\ln f_{p}^{-}+f_{p}^{+}\ln f_{p}^{+}\right).$ (21)
The grand potential is therefore given by
$\Omega^{\text{HF}}$$\equiv$ $U^{\text{HF}}-\mu^{\text{HF}}\bar{N}-TS^{\text{HF}}$ (23)
$=$ $\sum_{p}\left(\epsilon_{p}^{\text{HF}}-\mu^{\text{HF}}\right)f_{p}^{-}-\frac{1}{2}\sum_{p,q}\langle pq||pq\rangle f_{p}^{-}f_{q}^{-}$
$+\frac{1}{\beta}\sum_{p}\left(f_{p}^{-}\ln f_{p}^{-}+f_{p}^{+}\ln f_{p}^{+}\right).$
These thermodynamic functions together satisfyGu and Hirata (2024) the thermodynamic relations such as Eq. (23) by construction as well as
$-\frac{\partial\Omega^{\text{HF}}}{\partial\mu^{\text{HF}}}$ $=$ $\bar{N},$ (24)
$-\frac{\partial\Omega^{\text{HF}}}{\partial T}$ $=$ $S^{\text{HF}},$ (25)
$\frac{\partial\Omega^{\text{HF}}}{\partial f_{p}^{-}}$ $=$ $0$ (26)
The last identity proves that $\Omega^{\text{HF}}$ is indeed minimizedMermin (1963); Gu and Hirata (2024) by the orbitals that diagonalize $\bm{\epsilon}^{\text{HF}}$ [Eq. (17)] and by the one-particle density matrix whose eigenvalues are the Fermi–Dirac distribution functions of Eq. (18).
IV Thermal quasi-particle theory
Second-order thermal quasi-particle theory [thermal QP(2) theory] is postulated by its internal energy,
$U^{\text{QP(2)}}$$\equiv$ $\langle E^{(0)}\rangle+\langle E^{(1)}\rangle+\langle E^{(2)}\rangle.$ (27)
The last term $\langle E^{(2)}\rangle$ is the thermal average of second-order (degenerate) perturbation energiesHirschfelder and Certain (1974) and has been reduced in Eqs. (C7) and (C8) of Ref. Hirata and Jha, 2020 to the following sum-over-orbitals formula:
$\langle E^{(2)}\rangle$ $=$ $\sum_{p,q}^{\text{denom.}\neq 0}\frac{F_{qp}F_{pq}}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{p}^{-}f_{q}^{+}$
$+\frac{1}{4}\sum_{p,q,r,s}^{\text{denom.}\neq 0}\frac{\langle pq||rs\rangle\langle rs||pq\rangle}{\epsilon_{p}^{(0)}+\epsilon_{q}^{(0)}-\epsilon_{r}^{(0)}-\epsilon_{s}^{(0)}}f_{p}^{-}f_{q}^{-}f_{r}^{+}f_{s}^{+}$
with
$F_{pq}\equiv\epsilon_{pq}^{\text{HF}}-\delta_{pq}\epsilon_{p}^{(0)},$ (29)
where “denom.$\neq 0$” restricts the summation to over those indexes whose denominator is nonzero.Hirata and Jha (2020); Hirata (2021a)
It is a natural finite-temperature generalization of second-order many-body perturbation energy,Bartlett (1981); Shavitt and Bartlett (2009) but is considerably simpler than the full second-order correction to the internal energy $U^{(2)}$ of finite-temperature MBPT [Eq. (74) of Ref. Hirata and Jha, 2020]. The simplification is at least partly justified by the equally simpler treatments of the chemical-potential and entropy contributions to adhere to the quasi-independent-particle framework that the theory adopts. The first term of $\langle E^{(2)}\rangle$ [Eq. (LABEL:E2)] is the so-called non-HF term (or more precisely, non-thermal-HF term, in this case),Shavitt and Bartlett (2009) which is zero when thermal HF theory is used as the reference (where $F_{pq}=0$). The diagrammatic representationHirata (2021a) of $\langle E^{(2)}\rangle$ is given in Fig. 1.
Figure 1: Shavitt–Bartlett-style diagramsShavitt and Bartlett (2009); Hirata (2021a) of $\langle E^{(2)}\rangle$ of Eq. (LABEL:E2). Each downgoing (upgoing) edge represents $f_{p}^{-}$ ($f_{p}^{+}$), a dashed line is an anti-symmetrized two-electron integral, a dashed line with a circle denotes $F_{pq}$, and a wiggly (resolvent) line is a denominator factor.
We thus have
$U^{\text{QP(2)}}$$\equiv$ $\sum_{p}\epsilon_{p}^{\text{HF}}f_{p}^{-}-\frac{1}{2}\sum_{p,q}\langle pq||pq\rangle f_{p}^{-}f_{q}^{-}$
$+\sum_{p,q}^{\text{denom.}\neq 0}\frac{F_{qp}F_{pq}}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{p}^{-}f_{q}^{+}$
$+\frac{1}{4}\sum_{p,q,r,s}^{\text{denom.}\neq 0}\frac{\langle pq||rs\rangle\langle rs||pq\rangle}{\epsilon_{p}^{(0)}+\epsilon_{q}^{(0)}-\epsilon_{r}^{(0)}-\epsilon_{s}^{(0)}}f_{p}^{-}f_{q}^{-}f_{r}^{+}f_{s}^{+}.$
The entropy formula retains the one-particle picture and takes the form,
$S^{\text{QP(2)}}$$\equiv$ $-k_{\text{B}}\sum_{p}\left(f_{p}^{-}\ln f_{p}^{-}+f_{p}^{+}\ln f_{p}^{+}\right),$ (31)
which differs materially from the $S^{(0)}+S^{(1)}+S^{(2)}$ of Refs. Hirata and Jha, 2020; Hirata, 2021a, but is more in line with thermal HF or Fermi–Dirac theory. The corresponding grand potential is then defined by
$\Omega^{\text{QP(2)}}$$\equiv$ $U^{\text{QP(2)}}-\mu^{\text{QP(2)}}\bar{N}-TS^{\text{QP(2)}}$ (33)
$=$ $\sum_{p}\left(\epsilon_{p}^{\text{HF}}-\mu^{\text{QP(2)}}\right)f_{p}^{-}-\frac{1}{2}\sum_{p,q}\langle pq||pq\rangle f_{p}^{-}f_{q}^{-}$
$+\frac{1}{\beta}\sum_{p}\left(f_{p}^{-}\ln f_{p}^{-}+f_{p}^{+}\ln f_{p}^{+}\right)+\langle E^{(2)}\rangle,$
where $\mu^{\text{QP(2)}}$ is the chemical potential, which is determined by the same electroneutrality condition of Fermi–Dirac theory, i.e.,
$\bar{N}$ $=$ $\sum_{p}f_{p}^{-}.$ (34)
Spinorbitals labeled by $p$ and $q$ are those of a reference theory, which is typically but not limited to zero-temperature HF theory or thermal HF theory at the same $T$. No orbital rotation by matrix diagonalization is performed in this ansatz, but $\Omega^{\text{QP(2)}}$ is still minimized with respect to the diagonal elements of the one-electron density matrix, which are none other than the $f_{p}^{-}$.
These $f_{p}^{\mp}$ take the form of the Fermi–Dirac distribution functions, when $\Omega^{\text{QP(2)}}$ is minimized by varying $f_{p}^{-}$ (see below for a proof).Gu and Hirata (2024) They are defined with the second-order quasi-particle energy $\epsilon_{p}^{\text{QP(2)}}$ and chemical potential $\mu^{\text{QP(2)}}$, i.e.,
$f_{p}^{-}$ $=$ $\frac{1}{1+e^{\beta(\epsilon_{p}^{\text{QP(2)}}-\mu^{\text{QP(2)}})}},$ (35)
$f_{p}^{+}$ $=$ $1-f_{p}^{-}=\frac{e^{\beta(\epsilon_{p}^{\text{QP(2)}}-\mu^{\text{QP(2)}})}}{1+e^{\beta(\epsilon_{p}^{\text{QP(2)}}-\mu^{\text{QP(2)}})}}$ (36)
with $\epsilon_{p}^{\text{QP(2)}}$ given by
$\epsilon_{p}^{\text{QP(2)}}$ $=$ $\epsilon_{p}^{\text{HF}}+\Sigma_{pp}^{(2)}.$ (37)
Here, $\Sigma_{pp}^{(2)}$ is a finite-temperature analogue of the second-order self-energy of MBGF.Hirata et al. (2017) In a zero-temperature reference, where $\epsilon_{p}^{(0)}$ is constant of $T$, it takes the form of
$\Sigma_{pp}^{(2)}$ $=$ $\sum_{q}^{\text{denom.}\neq 0}\frac{F_{qp}F_{pq}}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{q}^{+}-\sum_{q}^{\text{denom.}\neq 0}\frac{F_{qp}F_{pq}}{\epsilon_{q}^{(0)}-\epsilon_{p}^{(0)}}f_{q}^{-}$ (38)
$+\sum_{q,r}^{\text{denom.}\neq 0}\frac{\langle qr||pr\rangle F_{pq}}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{p}^{-}f_{q}^{+}$
$+\sum_{q,r}^{\text{denom.}\neq 0}\frac{F_{qp}\langle pr||qr\rangle}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{p}^{-}f_{q}^{+}$
$+\frac{1}{2}\sum_{q,r,s}^{\text{denom.}\neq 0}\frac{\langle pq||rs\rangle\langle rs||pq\rangle}{\epsilon_{p}^{(0)}+\epsilon_{q}^{(0)}-\epsilon_{r}^{(0)}-\epsilon_{s}^{(0)}}f_{q}^{-}f_{r}^{+}f_{s}^{+}$
$-\frac{1}{2}\sum_{q,r,s}^{\text{denom.}\neq 0}\frac{\langle rs||pq\rangle\langle pq||rs\rangle}{\epsilon_{r}^{(0)}+\epsilon_{s}^{(0)}-\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{q}^{+}f_{r}^{-}f_{s}^{-},$
which is obtained by demanding all thermodynamic relations be obeyed by thermodynamic functions defined by this theory (see below for a derivation). A diagrammatic representation of $\Sigma_{pp}^{(2)}$ is given in Fig. 2.
[…]
The third and fourth terms (diagrams) are non-HF termsShavitt and Bartlett (2009) and are also related to the energy-independent diagrams.Purvis and Öhrn (1975); Schirmer and Angonoa (1989)
When a zero-temperature HF reference is used, in the $T=0$ limit, the first four terms vanish because $F_{pq}=0$, leaving only the last two terms as the zero-temperature second-order self-energy in the diagonal, frequency-independent approximation. Their diagrams (the last two in Fig. 2) are, therefore, isomorphic to the two second-order self-energy diagrams (e.g., Fig. 4 of Ref. Hirata et al., 2017), although their scope is different (the former are for $T\geq 0$, while the latter are only for $T=0$).
Figure 2: Second-order thermal Dyson self-energy diagrams in the zero-temperature reference [Eq. (38)], which are obtained by cutting an edge in each of the internal energy diagrams in Fig. 1 and trimming the dangling lines. See also the caption of Fig. 1 as well as Refs. Hirata, 2021a; Hirata et al., 2017.
With these definitions, the thermodynamic relation of Eq. (33) is satisfied by construction. The next one,
$-\frac{\partial\Omega^{\text{QP(2)}}}{\partial\mu^{\text{QP(2)}}}$ $=$ $\bar{N},$ (39)
is also obeyed, which can be confirmed by explicit evaluation of the derivative.
$-\frac{\partial\Omega^{\text{QP(2)}}}{\partial\mu^{\text{QP(2)}}}$ $=$ $-\sum_{p}\frac{\partial\epsilon_{p}^{\text{HF}}}{\partial\mu^{\text{QP(2)}}}f_{p}^{-}+\sum_{p}f_{p}^{-}$ (40)
$-\sum_{p}\left(\epsilon_{p}^{\text{HF}}-\mu^{\text{QP(2)}}\right)\frac{\partial f_{p}^{-}}{\partial\mu^{\text{QP(2)}}}$
$+\sum_{p,q}\langle pq||pq\rangle\frac{\partial f_{p}^{-}}{\partial\mu^{\text{QP(2)}}}f_{q}^{-}$
$-\frac{1}{\beta}\sum_{p}\frac{\partial f_{p}^{-}}{\partial\mu^{\text{QP(2)}}}\ln f_{p}^{-}-\frac{1}{\beta}\sum_{p}\frac{\partial f_{p}^{-}}{\partial\mu^{\text{QP(2)}}}$
$-\frac{1}{\beta}\sum_{p}\frac{\partial f_{p}^{+}}{\partial\mu^{\text{QP(2)}}}\ln f_{p}^{+}-\frac{1}{\beta}\sum_{p}\frac{\partial f_{p}^{+}}{\partial\mu^{\text{QP(2)}}}-\frac{\partial\langle E^{(2)}\rangle}{\partial\mu^{\text{QP(2)}}}$
$=$ $\sum_{p}f_{p}^{-}-\sum_{p}\left(\epsilon_{p}^{\text{HF}}-\mu^{\text{QP(2)}}\right)\frac{\partial f_{p}^{-}}{\partial\mu^{\text{QP(2)}}}$
$-\frac{1}{\beta}\sum_{p}\frac{\partial f_{p}^{-}}{\partial\mu^{\text{QP(2)}}}\ln\frac{f_{p}^{-}}{f_{p}^{+}}-\frac{\partial\langle E^{(2)}\rangle}{\partial\mu^{\text{QP(2)}}}$
$=$ $\sum_{p}f_{p}^{-}+\sum_{p}\Sigma_{pp}^{(2)}\frac{\partial f_{p}^{-}}{\partial\mu^{\text{QP(2)}}}-\frac{\partial\langle E^{(2)}\rangle}{\partial\mu^{\text{QP(2)}}}$
$=$ $\sum_{p}f_{p}^{-}=\bar{N},$
where we used
$\frac{\partial f_{p}^{+}}{\partial\mu^{\text{QP(2)}}}$ $=$ $-\frac{\partial f_{p}^{-}}{\partial\mu^{\text{QP(2)}}},$ (41)
$\ln\frac{f_{p}^{-}}{f_{p}^{+}}$ $=$ $-\beta\left(\epsilon_{p}^{\text{QP(2)}}-\mu^{\text{QP(2)}}\right),$ (42)
but not the explicit derivative $\partial f_{p}^{-}/\partial\mu^{\text{QP(2)}}$, which is a solution of a system of linear equations and cannot be written in a closed form.Gu and Hirata (2024) This justifies Eq. (34).
In the penultimate equality of Eq. (40), we also used
$\sum_{p}\Sigma_{pp}^{(2)}\frac{\partial f_{p}^{-}}{\partial\mu^{\text{QP(2)}}}=\frac{\partial\langle E^{(2)}\rangle}{\partial\mu^{\text{QP(2)}}},$ (43)
which can be verified by using Eqs. (LABEL:E2) and (38). In fact, the form of $\Sigma_{pp}^{(2)}$ is determined in the first place from the general condition,
$\Sigma_{pp}^{(n)}\equiv\frac{\partial\langle E^{(n)}\rangle}{\partial f_{p}^{-}},$ (44)
which ensures that Eq. (43) and all similar relations be satisfied. The $\Sigma_{pp}^{(2)}$ formula of Eq. (38) was postulated by substituting Eq. (LABEL:E2) into the above formula.
Equation (44) is a finite-temperature generalizationMatsubara (1955); Luttinger and Ward (1960) of the $n$th-order self-energy of MBGF,Hirata et al. (2017, 2015) which is diagonal and frequency independent. Owing to the diagonal construction, the second- and higher-order thermal quasi-particle theory does not involve rotation of orbitals, unlike thermal HF theoryMermin (1963); Gu and Hirata (2024) or Feynman–Dyson MBGF,Hirata et al. (2017) which define the HF or Dyson orbitalsOrtiz (2020) as those that bring the Fock or self-energy matrix into a diagonal form. Despite the absence of orbital rotation, $\Omega^{\text{QP(2)}}$ is still minimized by varying $f_{p}^{-}$ (see below).
Diagrammatically, the differentiation with respect to $f_{p}^{-}$ corresponds to opening a closed, internal-energy diagram by deleting an edge.Hirata et al. (2017) A possible physical meaning of thermal self-energies or thermal quasi-particle energies, which are temperature dependent, is given and discussed in Sec. VI.
The following thermodynamic relation,
$-\frac{\partial\Omega^{\text{QP(2)}}}{\partial T}$ $=$ $S^{\text{QP(2)}},$ (45)
is also obeyed. This too can be confirmed by explicit differentiation.
$-\frac{\partial\Omega^{\text{QP(2)}}}{\partial T}$ $=$ $k_{\text{B}}\beta^{2}\frac{\partial\Omega^{\text{QP(2)}}}{\partial\beta}$ (46)
$=$ $k_{\text{B}}\beta^{2}\sum_{p}\frac{\partial\epsilon_{p}^{\text{HF}}}{\partial\beta}f_{p}^{-}$
$+k_{\text{B}}\beta^{2}\sum_{p}\left(\epsilon_{p}^{\text{HF}}-\mu^{\text{QP(2)}}\right)\frac{\partial f_{p}^{-}}{\partial\beta}$
$-k_{\text{B}}\beta^{2}\sum_{p,q}\langle pq||pq\rangle\frac{\partial f_{p}^{-}}{\partial\beta}f_{q}^{-}$
$-k_{\text{B}}\left(f_{p}^{-}\ln f_{p}^{-}+f_{p}^{+}\ln f_{p}^{+}\right)$
$+k_{\text{B}}\beta\sum_{p}\frac{\partial f_{p}^{-}}{\partial\beta}\ln\frac{f_{p}^{-}}{f_{p}^{+}}$
$+k_{\text{B}}\beta^{2}\frac{\partial\langle E^{(2)}\rangle}{\partial\beta}$
$=$ $-k_{\text{B}}\left(f_{p}^{-}\ln f_{p}^{-}+f_{p}^{+}\ln f_{p}^{+}\right)$
$-k_{\text{B}}\beta^{2}\sum_{p}\frac{\partial\Sigma_{pp}^{(2)}}{\partial\beta}f_{p}^{-}+k_{\text{B}}\beta^{2}\frac{\partial\langle E^{(2)}\rangle}{\partial\beta}$
$=$ $-k_{\text{B}}\left(f_{p}^{-}\ln f_{p}^{-}+f_{p}^{+}\ln f_{p}^{+}\right)=S^{\text{QP(2)}},$
where we used Eqs. (42) and (44), but not the explicit derivative $\partial f_{p}^{-}/\partial\beta$.Gu and Hirata (2024) This justifies Eq. (31).
It can also be shown that thermal QP(2) theory is variational with respect to the diagonal elements of its one-electron density matrix, but not with respect to orbitals.
$\frac{\partial\Omega^{\text{QP(2)}}}{\partial f_{p}^{-}}$ $=$ $\left(\epsilon_{p}^{\text{HF}}-\mu^{\text{QP(2)}}\right)+\frac{1}{\beta}\ln\frac{f_{p}^{-}}{f_{p}^{+}}+\frac{\partial\langle E^{(2)}\rangle}{\partial f_{p}^{-}}$
$=$ $\left(\epsilon_{p}^{\text{HF}}-\mu^{\text{QP(2)}}\right)-\left(\epsilon_{p}^{\text{QP(2)}}-\mu^{\text{QP(2)}}\right)+\Sigma_{pp}^{(2)}=0,$
where Eq. (44) was used. This proves that the $f_{p}^{-}$ functions that minimize $\Omega^{\text{QP(2)}}$ take the forms of the Fermi–Dirac distribution functions, justifying Eqs. (35) and (36).
Insofar as Eq. (44) is satisfied (and thus all thermodynamic relations are obeyed and the variationality is ensured), there is considerable latitude in selecting the functional form of an approximate internal energy. For instance, the most notable departure of $U^{\text{QP(2)}}$ in Eq. (LABEL:QP:U) from the lengthy, but full $U^{(2)}$ of Eq. (74) of Ref. Hirata and Jha, 2020 is the absence of the so-called anomalous-diagram terms,Kohn and Luttinger (1960); Luttinger and Ward (1960) which sum over the set of indexes whose fictitious denominator is zero.Hirata and Jha (2020); Hirata (2021a) They are purposefully neglected in our ansatz because they are found to cause a severe Kohn–Luttinger nonconvergence problem in $\Sigma_{pp}^{(2)}$,Kohn and Luttinger (1960); Luttinger and Ward (1960); Hirata (2021b, 2022) and are undesirable from both theoretical and practical points of view. This issue will be expounded in Appendix A.
When the reference wave function is supplied by a finite-temperature theory, its $\epsilon_{p}^{(0)}$ depends on temperature and the formulas for the thermal self-energies need to be adjusted accordingly. In the thermal HF reference, the second-order thermal self-energy $\Sigma_{pp}^{(2)}$ is found to contain third-order energy-independent diagramsPurvis and Öhrn (1975); Schirmer and Angonoa (1989) among others (not considered further in this study).
Not only does finite-temperature MBPT suffer from the aforementioned Kohn–Luttinger nonconvergence,Hirata (2021b, 2022) but the Feynman–Dyson perturbation series of zero-temperature MBGF is also shown to display even severer divergence for most low- and high-lying states.Hirata et al. (2024) The impact of the latter divergence on thermal quasi-particle theories will be touched upon in Appendix B.
V Thermodynamic functions
Table 1: Grand potential $\Omega$ (in $E_{\text{h}}$) of the hydrogen fluoride molecule (0.9168 Å) in the STO-3G basis set as a function of temperature ($T$).
$T$ / K QP(1)=HF [^1] First-order thermal quasi-particle theory or thermal HF theory. MBPT(2) [^2] Second-order finite-temperature MBPTHirata and Jha (2019, 2020); Hirata (2021a) with zero-temperature HF reference. QP(2) [^3] Second-order thermal quasi-particle theory with zero-temperature HF reference. FCI [^4] Thermal FCI theory.Kou and Hirata (2014)
$10^{4}$ $-99.50758$ $-99.94001$ $-99.94179$ $-99.94377$
$10^{5}$ $-101.02137$ $-103.48646$ $-101.30273$ $-102.10659$
$10^{6}$ $-150.56294$ $-151.43748$ $-150.56368$ $-151.24440$
$10^{7}$ $-729.93806$ $-730.10421$ $-729.93862$ $-730.09519$
$10^{8}$ $-6846.98049$ $-6847.00261$ $-6846.98055$ $-6847.00247$
Figure 3: The deviation from thermal FCI theory in grand potential $\Omega$ of the hydrogen fluoride molecule (0.9168 Å) in the STO-3G basis set as a function of temperature. The “HF” stands for thermal HF theory [thermal QP(1) theory], “MBPT(2)” denotes the second-order finite-temperature many-body perturbation theory with the zero-temperature HF reference, “QP(2)” designates the second-order thermal quasi-particle theory with the zero-temperature HF reference.
Table 2: Same as Table 1 but for internal energy $U$ (in $E_{\text{h}}$).
[…]
$10^{4}$ $-98.57076$ $-98.58809$ $-98.58810$ $-98.59658$
$10^{5}$ $-97.94385$ $-97.86604$ $-97.94314$ $-98.04938$
$10^{6}$ $-96.79410$ $-96.99284$ $-96.78681$ $-96.94534$
$10^{7}$ $-92.02773$ $-92.05724$ $-92.02752$ $-92.05557$
$10^{8}$ $-88.48266$ $-88.48744$ $-88.48268$ $-88.48740$
Figure 4: Same as Fig. 3 but for internal energy $U$.
Table 3: Same as Table 1 but for chemical potential $\mu$ (in $E_{\text{h}}$).
[…]
$10^{4}$ $0.09368$ $0.13519$ $0.13537$ $0.13472$
$10^{5}$ $0.20722$ $0.42903$ $0.23324$ $0.29568$
$10^{6}$ $3.80022$ $3.87744$ $3.79999$ $3.85990$
$10^{7}$ $46.85490$ $46.86975$ $46.85489$ $46.86892$
$10^{8}$ $504.65280$ $504.65478$ $504.65280$ $504.65476$
Figure 5: Same as Fig. 3 but for chemical potential $\mu$.
Table 4: Same as Table 1 but for entropy $S$ (in $k_{\text{B}}$).
[…]
$10^{4}$ $0.00000$ $0.00001$ $0.00001$ $0.00011$
$10^{5}$ $3.17451$ $4.20017$ $3.24367$ $3.47472$
$10^{6}$ $4.97871$ $4.94828$ $4.98197$ $4.95769$
$10^{7}$ $5.34800$ $5.34763$ $5.34803$ $5.34766$
$10^{8}$ $5.40597$ $5.40596$ $5.40597$ $5.40596$
Figure 6: Same as Fig. 3 but for entropy $S$.
Tables 14 list the thermodynamic functions ($\Omega$, $U$, $\mu$, and $S$) of thermal HF [thermal QP(1)],Gu and Hirata (2024) thermal QP(2), fintie-temperature MBPT(2),Hirata and Jha (2020); Hirata (2021a) and thermal FCI theoriesKou and Hirata (2014) for an ideal gas of the identical hydrogen fluoride molecules in a wide range of temperature ($T$). Figures 36 plot the deviations from thermal FCI benchmarks, the latter being exact within a basis set. The thermal QP(2) and finite-temperature MBPT(2) calculations were based on the zero-temperature HF reference. See Ref. Gu and Hirata, 2024 for the data for the same system from Fermi–Dirac [thermal QP(0)] theory and other thermal mean-field theories.
In all cases, thermal QP(2) results are close to finite-temperature MBPT(2) ones at low $T$ and are more accurate than thermal HF theory. This is expected because both thermal QP(2) theory and finite-temperature MBPT(2) reduce to zero-temperature MBPT(2), accounting for electron correlation, while thermal HF theory and its zero-temperature counterpart do not. At higher $T$, thermal QP(2) theory quickly approaches thermal HF theory, where their quasi-independent-particle picture becomes relatively less erroneous. However, at $T=10^{6}$ K and $10^{7}$ K, finite-temperature MBPT(2) is systematically more accurate as it accounts for the effect of electron correlation in $S$ and $\mu$ and it uses more complete second-order corrections (including anomalous-diagram terms) in $\Omega$ and $U$. At the intermediate temperature of $T=10^{5}$ K, thermal QP(2) theory (or even thermal HF theory) outperforms finite-temperature MBPT(2). This may be ascribed to the fact that thermal QP(2) theory includes the effect of electron correlation in the orbital energies, and thus higher-order correlation corrections in a manner similar in spirit to self-consistent Green’s function methods,Luttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Van Neck et al. (1991); Dickhoff (1999); Dickhoff and Barbieri (2004); Dahlen and van Leeuwen (2005); Barbieri (2006); Barbieri and Dickhoff (2009); Phillips and Zgid (2014); Neuhauser et al. (2017); Tarantino et al. (2017); Kadanoff and Baym (2018); Coveney and Tew (2023) which are said to work for strong correlation. However, this similarity must not be overstated because self-consistent Green’s function methods replace reference orbitals and orbital energies by electron-correlated counterparts, while thermal QP(2) theory uses correlated orbital energies only in the Fermi–Dirac distribution functions and thus reduces identically to MBPT(2) at $T=0$.
VI Thermal quasi-particle energies
Equation (44) can be rewritten as
$\epsilon_{p}\equiv\frac{\partial U}{\partial f_{p}^{-}},$ (48)
encompassing Fermi–Dirac, thermal HF, and thermal QP(2) theories.
The $\epsilon_{p}$ has the literal physical meaning of the increase in the internal energy upon infusion of an infinitesimal fraction of an electron in the $p$th spinorbital. While it is not irrational to speak about a fraction of an electron here because the process in question is the thermal average of the same processes involving an infinite number of particles, it is unclear how this quantity can be used in relation to observables. The $\epsilon_{p}$ at $T=0$ is rigorously connected with an ionization or electron-attachment energy via Koopmans’ theoremSzabo and Ostlund (1982) or MBGF.Hirata et al. (2017) However, such an interpretation will face difficulty when $T\gg 0$ because ionization or electron attachment at finite temperature will involve different orbitals at different probabilities, even in a strict independent-particle picture.
Just for convenience of the following analysis, let us define the internal energy $U(N)$ as a function of the average number of electrons $N$ (which may now differ from $\bar{N}$) by the respective formula of $U$ with only the chemical potential adjusted. For instance, the internal energy of thermal FCI theory is given by
$U^{\text{FCI}}(N)$ $=$ $\frac{\sum_{I}E_{I}^{\text{FCI}}e^{-\beta(E_{I}^{\text{FCI}}-\mu N_{I})}}{\sum_{I}e^{-\beta(E_{I}^{\text{FCI}}-\mu N_{I})}},$ (49)
in which only the $\mu$ is adjusted for a given $N$ to satisfy,
$N$ $=$ $\frac{\sum_{I}N_{I}e^{-\beta(E_{I}^{\text{FCI}}-\mu N_{I})}}{\sum_{I}e^{-\beta(E_{I}^{\text{FCI}}-\mu N_{I})}}.$ (50)
This is a valid and in fact, exact (within a basis set) treatment of thermodynamics if the particles are not electrically charged. For electrons, however, it describes a physically unrealistic, massively charged plasma, whose energy is not even extensive.Fisher and Ruelle (1966); Dyson and Lenard (1967); Hirata and Ohnishi (2012) Nevertheless, $U(N)$ is computationally well defined even for electrons thanks to the ideal-gas assumption (i.e., no interactions between molecules) and will be useful for the following analysis.
For Fermi–Dirac, thermal HF, and thermal QP(2) theories, the internal energy $U(N)$ as a function of $N$ is defined by Eqs. (1), (13), and (LABEL:QP:U), respectively, with its $\mu$ determined for a given $N$ by
$N=\sum_{p}f_{p}^{-},$ (51)
which no longer stands for electroneutrality. Then, $U(N-1)-U(N)$ may be considered as a crude estimate of the thermal average of the first ionization energies. Likewise, $U(N)-U(N+1)$ is an approximate thermal first electron-attachment energy. At $T=0$, they correspond to the energy (the sign reversed) of the highest-occupied and lowest-unoccupied molecular or DysonOrtiz (2020) orbitals (HOMO and LUMO), respectively, according to Koopmans’ theoremSzabo and Ostlund (1982) or MBGF.Hirata et al. (2017) Since $U(N)$ is defined and computed with $\epsilon_{p}$, these internal-energy differences may furnish a physical meaning of the latter, which is indirect, yet potentially useful in relation to observables.
We call $U(N)-U(N-1)$ the thermal ionization energy and $U(N+1)-U(N)$ the thermal electron-attachment energy. Note the uncommon sign convention.
The $U(N)$ thus defined are computed for an ideal gas of identical hydrogen fluoride molecules by thermal FCI, thermal QP(2), and thermal HF theories in Figs. 7, 8, and 9, respectively. The thermal ionization energies $U(N)-U(N-1)$ of these three methods are compared with one another and also with the thermal HOMO energies in Table 5. The same data for electron-attachment energies and LUMO are compiled in Table 6.
Figure 7: Internal energy of thermal FCI theory $U_{\text{FCI}}$ (in $E_{\text{h}}$) of the hydrogen fluoride molecule (0.9168 Å) in the STO-3G basis set as a function of the average number of electrons $N$ relative to the value for $N=\bar{N}=10$.
Figure 8: Same as Fig. 7 but for the internal energy of thermal QP(2) theory $U_{\text{QP(2)}}$.
Figure 9: Same as Fig. 7 but for the internal energy of thermal HF theory $U_{\text{HF}}$.
Table 5: Thermal ionization energy (in $E_{\text{h}}$) of the hydrogen fluoride molecule (0.9168 Å) in the STO-3G basis set as a function of temperature ($T$).
Thermal HOMO energy $U(N)-U(N-1)$
$T$ / K QP(1)=HF [^1] First-order thermal quasi-particle theory or thermal HF theory. QP(2) [^2] Second-order thermal quasi-particle theory with zero-temperature HF reference. QP(1)=HF [^1] First-order thermal quasi-particle theory or thermal HF theory. QP(2) [^2] Second-order thermal quasi-particle theory with zero-temperature HF reference. FCI [^3] Thermal FCI theory.Kou and Hirata (2014)
00444Ionization energy from the corresponding MBGF.Hirata et al. (2017, 2024) $-0.46417$ $-0.39557$ $-0.46417$ $-0.39557$ $-0.40429$
$10^{4}$ $-0.46417$ $-0.39557$ $-0.46589$ $-0.39603$ $-0.40468$
$10^{5}$ $-0.45147$ $-0.46765$ $-0.21004$ $-0.18425$ $-0.32041$
$10^{6}$ $-0.57384$ $-0.58577$ $-0.57181$ $-0.57425$ $-0.77028$
$10^{7}$ $-0.69361$ $-0.69692$ $-3.40146$ $-3.40165$ $-3.65153$
$10^{8}$ $-0.76988$ $-0.77209$ $-4.95141$ $-4.95143$ $-5.34456$
Table 6: Thermal electron-attachment energy (in $E_{\text{h}}$) of the hydrogen fluoride molecule (0.9168 Å) in the STO-3G basis set as a function of temperature ($T$).
Thermal LUMO energy $U(N+1)-U(N)$
$T$ / K QP(1)=HF [^1] First-order thermal quasi-particle theory or thermal HF theory. QP(2) [^2] Second-order thermal quasi-particle theory with zero-temperature HF reference. QP(1)=HF [^1] First-order thermal quasi-particle theory or thermal HF theory. QP(2) [^2] Second-order thermal quasi-particle theory with zero-temperature HF reference. FCI [^3] Thermal FCI theory.Kou and Hirata (2014)
00444Electron-attachment energy from the corresponding MBGF.Hirata et al. (2017, 2024) $0.62924$ $0.64424$ $0.62924$ $0.64424$ $0.65170$
$10^{4}$ $0.62924$ $0.64424$ $0.62924$ $0.64424$ $0.65170$
$10^{5}$ $0.48080$ $0.49392$ $0.07823$ $0.08932$ $0.40988$
$10^{6}$ $0.28118$ $0.20431$ $-0.55009$ $-0.55537$ $-0.12383$
$10^{7}$ $0.23384$ $0.13982$ $-3.14063$ $-3.14106$ $-2.71365$
$10^{8}$ $0.21118$ $0.11357$ $-4.90424$ $-4.90428$ $-4.48208$
From the tables, we observe that while the HOMO and LUMO energies of thermal quasi-particle theories correctly reduce to the ionization and electron-attachment energies of MBGF in the $T=0$ limit, they clearly have nothing to do with the approximate ionization or electron-attachment energies (as defined by $U(N)-U(N-1)$ or $U(N+1)-U(N)$ of thermal FCI theory) at $T\gg 0$, notwithstanding the questionable validity of the approximation. However, $U(N)-U(N-1)$ or $U(N+1)-U(N)$ calculated with $\epsilon_{p}$ of thermal quasi-particle theories are in line with the approximate ionization or electron-attachment energies of thermal FCI theory.
Thermal QP(2) theory is much closer to thermal FCI theory than thermal HF theory at low temperatures ($T\leq 10^{4}$ K), correctly accounting for electron correlation in thermal ionization and electron-attachment energies. However, at higher $T$, thermal HF and QP(2) theories are essentially the same, suggesting that the effect of electron correlation is surpassed by the lack thereof in the entropy and chemical potential in the quasi-independent-particle picture.
The figures reinforce these observations, but provide the following additional insights: As $T$ is lowered, the $U(N)$ curve begins to display the derivative discontinuity at $N=\bar{N}$, although the strict discontinuity does not occur until $T=0$. In both the low- and high-$T$ extremes, $U(N)$ has near-linear dependence on $N$ within each interval between adjacent integers, but at intermediate $T$, $U(N)$ is a convex curve with its minimum occurring away from $N=\bar{N}$.
Let us consider the slope of $U(N)$ at $N=\bar{N}$, which is another macroscopic quantity evaluable with $\epsilon_{p}$.
$\frac{\partial U}{\partial N}$ $=$ $\frac{\partial\mu}{\partial N}\sum_{p}\frac{\partial f_{p}^{-}}{\partial\mu}\frac{\partial U}{\partial f_{p}^{-}}$ (52)
$=$ $\frac{\sum_{p}f_{p}^{-}f_{p}^{+}\epsilon_{p}}{\sum_{p}f_{p}^{-}f_{p}^{+}},$
where we used Eq. (50) and
$\frac{\partial f_{p}^{-}}{\partial\mu}=\beta f_{p}^{-}f_{p}^{+}.$ (53)
This is the increase in the internal energy upon infusion of an infinitesimal fraction of an electron into every molecule in the ideal gas. It is a weighted average of $\epsilon_{p}$ over all spinorbitals with the $p$th spinorbital weight being $f_{p}^{-}f_{p}^{+}$. It corresponds neither to the ionization nor to the electron-attachment process because it is oblivious to the sign of the variation in $N$. However, in the $T=0$ limit, it distinguishes these two processes by exhibiting the derivative discontinuity, i.e.,
$\lim_{T\to 0}\left(\frac{\partial U}{\partial N}\right)_{N=\bar{N}^{-}}$ $=$ $\epsilon_{\text{HOMO}},$ (54)
$\lim_{T\to 0}\left(\frac{\partial U}{\partial N}\right)_{N=\bar{N}^{+}}$ $=$ $\epsilon_{\text{LUMO}}.$ (55)
Until the $T=0$ limit is strictly reached, however, the $U(N)$ curve is smooth everywhere and the slope approaches the midpoint of the HOMO and LUMO energies. For $T\approx 0$,
$\left(\frac{\partial U}{\partial N}\right)_{N=\bar{N}}$$\approx$ $\frac{\epsilon_{\text{HOMO}}+\epsilon_{\text{LUMO}}}{2},$ (56)
which is also equal to the $T=0$ limit of the $\mu$.Hirata and Jha (2019); Hirata (2021b)
Table 7: $\partial U/\partial N$ ((in $E_{\text{h}}$)) of of the hydrogen fluoride molecule (0.9168 Å) in the STO-3G basis set as a function of temperature ($T$).
$T$ / K QP(1)=HF [^1] First-order thermal quasi-particle theory or thermal HF theory [Eq. (52)]. QP(2) [^2] Second-order thermal quasi-particle theory with zero-temperature HF reference [Eq. (52)]. FCI [^3] Thermal FCI theory (numerical differentiation).Kou and Hirata (2014)
00444The average of ionization and electron-attachment energies of the corresponding MBGF.Hirata et al. (2017, 2024) $0.08253$ $0.12433$ $0.12371$
$10^{4}$ $0.08189$ $0.12461$ $0.12351$
$10^{5}$ $-0.07423$ $-0.05604$ $0.04959$
$10^{6}$ $-0.56092$ $-0.56475$ $-0.44097$
$10^{7}$ $-3.26523$ $-3.26554$ $-3.17327$
$10^{8}$ $-4.92771$ $-4.92775$ $-4.91206$
Table 7 compares the slope $\partial U/\partial N$ at $N=\bar{N}$ computed analytically by Eq. (52) for thermal HF and QP(2) theories against the slope of thermal FCI theory obtained by numerical differentiation. The three sets of values are in reasonable agreement with one another at all $T$ considered. At low $T$, thermal QP(2) theory is in much better agreement with thermal FCI theory than thermal HF theory as the former takes into account electron correlation. At higher $T$, thermal QP(2) and HF theories are more alike than thermal FCI theory, but they all seem to converge at the same high $T$ limit.
To summarize the results of the foregoing analysis, the orbital energies $\epsilon_{p}$ of thermal quasi-particle theories (including thermal HF theory) cannot be directly related to the ionization or electron-attachment energy from/into the $p$th orbital except at $T=0$. This is simply because an ionization in an ideal gas of molecules at $T\gg 0$ is not a removal of an electron from the $p$th orbital in every molecule; rather, it is a removal of an electron from various orbitals of all molecules at some probabilities (even in a strict independent-particle picture). This is, in turn, equivalent to a removal of some fraction of an electron from various orbitals of a single average molecule. The $\epsilon_{p}$ at $T>0$ signifies the increase in internal energy upon adding an infinitesimal fraction of an electron in the $p$th spinorbital of this “average” molecule. This quantity may not directly correspond to any observable, but it can be combined to produce a macroscopic thermodynamic observable, for which thermal QP(2) theory is in better agreement with thermal FCI theory than thermal HF theory.
VII Conclusions
The thermal quasi-particle theory has been introduced. It is a quasi-independent-particle theory and a finite-temperature extension of the widely used thermal HF theory. Its entropy and chemical-potential formulas are of the one-electron type,
$S\equiv-k_{\text{B}}\sum_{p}\left(f_{p}^{-}\ln f_{p}^{-}+f_{p}^{+}\ln f_{p}^{+}\right),$
and
$\bar{N}\equiv\sum_{p}f_{p}^{-}$
where the thermal population $f_{p}^{-}$ is also unchanged from the Fermi–Dirac distribution function,
$f_{p}^{-}\equiv\frac{1}{1+e^{\beta(\epsilon_{p}-\mu)}},$
and $f_{p}^{+}=1-f_{p}^{-}$. The orbital energies $\epsilon_{p}$ now become the ones that include the effect of electron correlation by being defined as the correlated, thermal one-electron energies in the spirit of MBGF,
$\epsilon_{p}\equiv\frac{\partial U}{\partial f_{p}^{-}},$
where $U$ is a correlated internal energy, whose mathematical form can be postulated on the basis of finite-temperature MBPT through any chosen order. When the grand potential is defined by
$\Omega\equiv U-\mu\bar{N}-TS,$
the theory satisfies all thermodynamic relations and the $\Omega$ is a variational minimum with respect to $f_{p}^{-}$.
Fermi–Dirac and thermal HF theories can be viewed as the zeroth- and first-order instances, respectively, in the hierarchy of thermal quasi-particle theories. It can be viewed as a diagonal, frequency-independent approximation to a finite-temperature extension of the inverse Dyson equation, which is an exact one-particle theory.Hirata et al. (2017) In other words, the $\epsilon_{p}$ thus defined is a finite-temperature extension of the Dyson self-energy in the diagonal, frequency-independent approximation. Generally, the thermal self-energy is shown to suffer from the severe kind of Kohn–Luttinger nonconvergence problem and so it is inappropriate to include anomalous-diagram terms in the definition of approximate $U$ or $\Omega$.
Our preliminary implementation and comparison with the thermal FCI benchmarks suggests that thermal QP(2) theory performs distinctly better than thermal HF theory at low $T$ and may even outperform finite-temperature MBPT(2) at intermediate $T$ by including electron-correlation effects in the orbital energies. At higher $T$, the entropy contribution dominates and the differences among various methods become relatively insignificant.
The direct physical meaning of $\epsilon_{p}$ is the increase in internal energy upon infusion of an infinitesimal fraction of an electron in the $p$th spinorbital. It does not correspond to the ionization energy or electron-attachment energy of a molecule at $T>0$. However, it can be used as ingredients from which electron-correlated thermodynamic observables can be computed.
Acknowledgements.
This work was supported by the U.S. Department of Energy (DoE), Office of Science, Office of Basic Energy Sciences under Grant No. DE-SC0006028 and also by the Center for Scalable Predictive methods for Excitations and Correlated phenomena (SPEC), which is funded by the U.S. DoE, Office of Science, Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences and Biosciences as part of the Computational Chemical Sciences (CCS) program at Pacific Northwest National Laboratory (PNNL) under FWP 70942. PNNL is a multi-program national laboratory operated by Battelle Memorial Institute for the U.S. DoE. The author is a Guggenheim Fellow of the John Simon Guggenheim Memorial Foundation.
Appendix A Kohn–Luttinger nonconvergence
Figure 10: Second-order internal energy diagrams of Eq. (59), which include anomalous (second and fourth) diagrams with no resolvent lines.
According to finite-temperature MBPT,Hirata and Jha (2019, 2020); Hirata (2021a) a more complete expression for the second-order internal energy should include the so-called anomalous-diagram contributions.Kohn and Luttinger (1960) It may read
$U^{\text{QP}}$$\equiv$ $\langle E^{(0)}\rangle+\langle E^{(1)}\rangle+U^{(2)}$ (57)
with
$U^{(2)}$$\equiv$ $\langle E^{(2)}\rangle-\beta\langle E^{(1)}E^{(1)}\rangle+\beta\langle E^{(1)}\rangle\langle E^{(1)}\rangle$ (58)
$=$ $\langle E^{(2)}\rangle-\beta\langle E^{(1)}E^{(1)}\rangle_{L},$
where subscript “$L$” means that only diagrammatically linked contributions are to be retained, and the last term (carrying a $\beta$ multiplier) is the anomalous-diagram term. This is still far from the whole second-order internal energy [Eq. (74) of Ref. Hirata and Jha, 2020], but the presence of the last anomalous-diagram term will make the following analysis meaningful.
The zeroth-order thermal averages in the above expression can be evaluatedHirata and Jha (2020); Hirata (2021a) as
$U^{(2)}$ $=$ $\sum_{p,q}^{\text{denom.}\neq 0}\frac{F_{qp}F_{pq}}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{p}^{-}f_{q}^{+}$ (59)
$-{\beta}\sum_{p,q}^{\text{denom.}=0}{F_{qp}F_{pq}f_{p}^{-}f_{q}^{+}}$
$+\frac{1}{4}\sum_{p,q,r,s}^{\text{denom.}\neq 0}\frac{\langle pq||rs\rangle\langle rs||pq\rangle}{\epsilon_{p}^{(0)}+\epsilon_{q}^{(0)}-\epsilon_{r}^{(0)}-\epsilon_{s}^{(0)}}f_{p}^{-}f_{q}^{-}f_{r}^{+}f_{s}^{+}$
$-\frac{\beta}{4}\sum_{p,q,r,s}^{\text{denom.}=0}{\langle pq||rs\rangle\langle rs||pq\rangle}f_{p}^{-}f_{q}^{-}f_{r}^{+}f_{s}^{+},$
whose diagrammatic representation is given in Fig. 10.
The second and fourth terms (carrying a $\beta$ multiplier but no denominator) are anomalous-diagram terms,Kohn and Luttinger (1960) causing the low-temperature breakdown known as the Kohn–Luttinger nonconvergence.Kohn and Luttinger (1960); Luttinger and Ward (1960); Hirata (2021b, 2022) Take the second term for example. The “denom.=0” restricts the summation to over those combinations of indexes whose fictitious denominator $\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}$ is zero. One such combination is $p=q$. As $T$ is lowered to zero (i.e., $\beta\to\infty$),Kohn and Luttinger (1960)
$\lim_{T\to 0}\beta f_{p}^{-}f_{p}^{+}=\delta\left(\mu-\epsilon_{p}^{(0)}\right),$ (60)
where the right-hand side is Dirac’s $\delta$ function, which is divergent when the chemical potential $\mu$ coincides with one of the orbital energies.Kohn and Luttinger (1960) Therefore, a Kohn–Luttinger nonconvergence occurs when $F_{pp}\neq 0$ and at the same time, the reference wave function is degenerate (i.e., the HOMO and LUMO have the same energy as $\mu$). While the divergence of this term can still be avoided by adopting a HF reference wherein $F_{pq}=0$,Luttinger and Ward (1960) the fourth term remains divergent for $p=r$ and $q=s$ in a degenerate reference.Hirata (2021b, 2022) It should be recalled that the second-order correction to the ground-state energy is always finite regardless of whether the zeroth-order ground state is degenerateHirschfelder and Certain (1974) or nondegenerate.Shavitt and Bartlett (2009) Hence, $U^{(2)}$ is not convergent at the correct zero-temperature limit, which is always finite.
Figure 11: Second-order thermal Dyson self-energy diagrams of Eq. (61), which include anomalous diagrams with no resolvent lines. They are obtained by cutting an edge in each of the internal energy diagrams of Fig. 10 and trimming the dangling lines.
As per Eq. (44), the corresponding thermal self-energy is obtained by taking the derivative of $U^{(2)}$ with respect to $f_{p}^{-}$, i.e.,
$\Sigma_{pp}^{(2)}$$\equiv$ $\frac{\partial U^{(2)}}{\partial f_{p}^{-}}$ (61)
$=$ $\sum_{q}^{\text{denom.}\neq 0}\frac{F_{qp}F_{pq}}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{q}^{+}-\sum_{q}^{\text{denom.}\neq 0}\frac{F_{qp}F_{pq}}{\epsilon_{q}^{(0)}-\epsilon_{p}^{(0)}}f_{q}^{-}$
$+\sum_{q,r}^{\text{denom.}\neq 0}\frac{\langle qr||pr\rangle F_{pq}}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{p}^{-}f_{q}^{+}$
$+\sum_{q,r}^{\text{denom.}\neq 0}\frac{F_{qp}\langle pr||qr\rangle}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{p}^{-}f_{q}^{+}$
$-{\beta}\sum_{q}^{\text{denom.}=0}F_{qp}F_{pq}f_{q}^{+}+{\beta}\sum_{q}^{\text{denom.}=0}F_{qp}F_{pq}f_{q}^{-}$
$-{\beta}\sum_{q,r}^{\text{denom.}=0}{\langle qr||pr\rangle F_{pq}f_{p}^{-}f_{q}^{+}}$
$-{\beta}\sum_{q,r}^{\text{denom.}=0}{F_{qp}\langle pr||qr\rangle f_{p}^{-}f_{q}^{+}}$
$+\frac{1}{2}\sum_{q,r,s}^{\text{denom.}\neq 0}\frac{\langle pq||rs\rangle\langle rs||pq\rangle}{\epsilon_{p}^{(0)}+\epsilon_{q}^{(0)}-\epsilon_{r}^{(0)}-\epsilon_{s}^{(0)}}f_{q}^{-}f_{r}^{+}f_{s}^{+}$
$-\frac{1}{2}\sum_{q,r,s}^{\text{denom.}\neq 0}\frac{\langle rs||pq\rangle\langle pq||rs\rangle}{\epsilon_{r}^{(0)}+\epsilon_{s}^{(0)}-\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{q}^{+}f_{r}^{-}f_{s}^{-}$
$-\frac{\beta}{2}\sum_{q,r,s}^{\text{denom.}=0}{\langle pq||rs\rangle\langle rs||pq\rangle}f_{q}^{-}f_{r}^{+}f_{s}^{+}$
$+\frac{\beta}{2}\sum_{q,r,s}^{\text{denom.}=0}{\langle rs||pq\rangle\langle pq||rs\rangle}f_{q}^{+}f_{r}^{-}f_{s}^{-}.$
The corresponding diagrams are shown in Fig. 11, which are obtained by opening each of the $U^{(2)}$ diagrams in Fig. 10 by deleting an edge.
This formula suffers from a more pervasive kind of the Kohn–Luttinger nonconvergence problem. Take the penultimate term for example. It sums over the $p=r$ and $q=s$ case, whose summand diverges as $T\to 0$ ($\beta\to\infty$) for any reference—degenerate or nondegenerate—because a $f_{p}^{-}$ factor is no longer there to create a $\delta$ function [Eq. (60)] which diverges only when $\mu=\epsilon_{p}^{(0)}$. In other words, a finite-temperature generalization of the self-energy, when it is derived more rigorously from finite-temperature MBPT and thus includes anomalous-diagram terms, is far more prone to divergence as $T\to 0$ and will be uselessly erroneous at low $T$. This is not surprising because in thermodynamics all states with any number of electrons are summed over in the partition function and they always include numerous exactly degenerate zeroth-order states.
This finding may bode ill for the Luttinger–Ward functionalLuttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Kozik et al. (2015); Rossi and Werner (2015); Welden et al. (2016); Gunnarsson et al. (2017); Lin and Lindsey (2018) because opening a diagram of this functional will result in a thermal self-energy or Green’s function diagram that may always suffer from the same severe low-temperature breakdown.
Appendix B Feynman–Dyson nonconvergence
In Ref. Hirata et al., 2024 was reported that a Feynman–Dyson perturbation expansion of the Dyson self-energy is always divergent in many frequency domains. Therefore, even apart from the divergence of the thermal self-energy caused by anomalous-diagram terms analyzed in Appendix A, the zero-temperature self-energy is expected to be divergent or at least excessively erroneous for low- and high-lying states. This is alleviated (or hidden from view) by the frequency-independent approximation inherent in the ansatz of thermal quasi-particle theory, but in a solid, whose self-energy forms continuous energy bands, we can no longer expect to be able to avoid this deep-rooted pathology.Hirata et al. (2024)
We therefore expect that in a solid-state implementation of thermal QP(2) theory (which is underway in our laboratory), we will introduce active orbitals (or active energy bands) and define and evaluate portions of $\langle E^{(2)}\rangle$ and of $\Sigma_{pp}^{(2)}$ involving these orbitals. These active orbitals are the ones whose zeroth-order energies fall within the so-called “central overlapping bracket” where the Feynman–Dyson perturbation expansion is guaranteed to have a nonzero radius of convergence.Hirata et al. (2024) Comparative performance of this method for solids will be presented in the future.